# A Note on Approximating Weighted Nash Social Welfare with Additive Valuations

### v1

@@ 2.3 The Analysis
Finally, we briefly discuss how to derandomize the rounding algorithm. We round the solution to the configuration LP in iterations, maintaining a fractional assignment $\bar{x}$ of items to agents; $\bar{x}=x$ initially. Let $\Delta$ be a large enough integer so that every $y_{i,S}$ is an integer multiply of $1/\Delta$. Focus on a fixed agent $i\in{\mathcal{A}}$ and consider the Nash Social Welfare instance containing $\Delta$ copies of $i$, and $\Delta\bar{x}_{ij}$ copies of each item $j\in{\mathcal{G}}$. Group the items as follows: the $\Delta$ most valuable items belong to the first group, the next $\Delta$ most valuable items belong to the second group, and so on. We define $\Phi_{i}$ to be the logarithm of the Nash Social Welfare of the worst allocation satisfying the following condition: every agent gets at most one item from each group. Fortunately, the worst allocation can be defined naturally: the first agent takes the most valuable item from each group, and the second agent takes the second most valuable item from each group, and so on. Thus $\Phi_{i}$ can be computed efficiently. We define $\Phi=\sum_{i}w_{i}\Phi_{i}$ to be the overall potential function. In the randomized version of the algorithm, one can define the rotation operation over the fractional matching between groups $G$ and items ${\mathcal{G}}$. In expectation the operation does not decrease $\Phi$. To derandomize the algorithm, we can perform the operation deterministically so that $\Phi$ does not decrease. The potential value $\Phi$ at the end of the algorithm is at least that at the beginning, which is at least the value of the configuration LP minus $1/e$. On the other hand, the logarithm of the Nash Social Welfare of the integral solution is exactly the final $\Phi$. Therefore, the Nash Social Welfare is at least $e^{-1/e}$ times the exponential of the value of the configuration LP.
@@ Appendix A Solving Configuration LP within an Additive Error of $\ln(1+\epsilon)$
Since (Conf-LP) has value strictly larger than $o$, the dual LP (4-6) is infeasible. We design an approximate separation oracle for the LP. Given some $\alpha\in{\mathbb{R}}_{\geq 0}^{\mathcal{G}}$ and $\beta\in{\mathbb{R}}^{\mathcal{A}}$ that does not satisfy (5), we can find some $i\in{\mathcal{A}}$ and $S\subseteq{\mathcal{G}}$ such that
$\sum_{j\in S}\alpha_{j}+\beta_{i}<w_{i}\ln\big{(}(1+\epsilon/2)v_{i}(S)\big{)}.$
The running time of the oracle is polynomial in the input size and $\frac{1}{\epsilon}$. This can be achieved using the standard dynamic programming technique: For a fixed $i\in{\mathcal{A}}$, to find the $S$, we guess the item $j^{*}\in S$ with the largest $v_{ij^{*}}$, coarsen the $v_{ij}$ values based on the guess, and run a dynamic programming to find the $S$.

### v2

@@ 2.3 The Analysis
It is easy to derandomize our algorithm, as we only need the constructed random matching between the groups $G$ and the items ${\mathcal{G}}$ to satisfy marginal probabilities. Therefore, we can decompose the fractional matching into a convex combination of polynomial number of matchings, and output the best matching in the combination.
@@ Appendix A Solving Configuration LP within an Additive Error of $\ln(1+\epsilon)$
Since (Conf-LP) has value strictly larger than $o$, the dual LP (4-6) is infeasible. We design an approximate separation oracle for the LP using approximate algorithms for the knapsack problems.
Given some $\alpha\in{\mathbb{R}}_{\geq 0}^{\mathcal{G}}$ and $\beta\in{\mathbb{R}}^{\mathcal{A}}$ that does not satisfy (5), we can find some $i\in{\mathcal{A}}$ and $S^{\prime}\subseteq{\mathcal{G}}$ such that
$\sum_{j\in S^{\prime}}\alpha_{j}+\beta_{i}<w_{i}\cdot\ln\big{(}(1+\epsilon/2)v_{i}(S^{\prime})\big{)}.$ (7)
The running time of the oracle is polynomial in the input size and $\frac{1}{\epsilon}$. This can be achieved using the standard dynamic programming technique. Suppose (5) is not satisfied for $i\in{\mathcal{A}}$ and $S\subseteq{\mathcal{G}}$. We can guess the agent $i$, and the item $j^{*}\in S$ with the largest $v_{ij^{*}}$. Then we discard the items $j$ with $v_{ij}>v_{ij^{*}}$, and round down $v_{ij}$ values for remaining items to the nearest integer multiple of $\frac{\epsilon\cdot v_{ij^{*}}}{2m}$. Let $\bar{v}_{ij}$ be the rounded value of item $j$. As $\bar{v}_{ij}$ values are integer multiples of $\frac{\epsilon\cdot v_{ij^{*}}}{2m}$ not exceeding $v_{ij^{*}}$, we can afford to guess $\bar{v}_{i}(S)$ as there are only $O(\frac{m^{2}}{\epsilon})$ possibilities for $\bar{v}_{i}(S)$, and use dynamic programming to find a set $S^{\prime}\subseteq{\mathcal{G}},S^{\prime}\ni j^{*}$ with $\bar{v}_{i}(S^{\prime})\geq\bar{v}_{i}(S)$ and $\sum_{j\in S^{\prime}}\alpha_{j}\leq\sum_{j\in S}\alpha_{j}$. So,
$v_{i}(S^{\prime})\geq\bar{v}_{i}(S^{\prime})\geq\bar{v}_{i}(S)\geq v_{i}(S)-m\cdot\frac{\epsilon\cdot v_{ij^{*}}}{2m}=v_{i}(S)-\frac{\epsilon\cdot v_{ij^{*}}}{2}\geq v_{i}(S)-\frac{\epsilon}{2}\cdot v_{i}(S^{\prime}).$
Therefore $v_{i}(S)\leq(1+\epsilon/2)v_{i}(S^{\prime})$. Then (7) follows from that (5) is not satisfied for $i,S$. The running time of the oracle is polynomial in the input size and $\frac{1}{\epsilon}$.